### v2

# Collective Creation of Intimacy: Exploring the Cosplay Commission Practice within the Otome Game Community in China

Abstract.
Cosplay commission (cos-commission) is a new form of commodified romantic companionship within the Otome game community in China. To explore the motivations, practices, experiences, and challenges, we conducted semi-structured interviews with 15 participants in different roles. Our findings reveal that cos-commission, as a hybrid activity, provides participants with a chance to collaboratively build meaningful connections. It also offers a pathway for personal exploration and emotional recovery. However, the vague boundary between performative roles and intimate interactions can give rise to unexpected negative outcomes, such as attachment-driven entanglements and post-commission ”withdrawal symptoms.” While digital platforms facilitate communication in cos-commissions, they often lack sufficient safeguards. This preliminary work provides insights into the formation process of hybrid intimate relationship and its potential to foster personalized, long-term support for mental well-being, and reveals potential privacy and security challenges.
video games, intimacy, virtual world, social media, online community
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.1122456journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Human-centered computing Empirical studies in HCI
Figure 1. Cosplay commission-related posts on different digital platforms (BiliBili, Douyin, Xiaohongshu and Xianyu).
[…]
Posts of cosplay commissions on different digital platforms.
1. Introduction
In recent years, Otome games (story-based romance video games for female players) have formed a large presence in the subcultural community of China. Although these games are played with the users spending their time and money to gain access to storylines and to form relationships with virtual characters, the offline community is even growing further, with more and more players and game developers hosting meetups in their areas. In tandem with this proclivity, a new form of commodified intimate experience has developed in the fandom, one that users have christened ”cosplay commission” or cos-commission in short. Through such practice, players can bring their virtual interactions into reality as they can pay cosplayers to physically represent the game characters and have a ”one-day romance” experience.
The rise of cos-commission has received attention from both academia (Lei et al., 2024; Ganzon, 2019; Wu et al., 2024) and the media (Wu, 2023; Zhu and Yang, 2024; Feng, 2018; Liang, 2019), reflecting the rapid nature of its growth. Online game communities have been one of the main areas of HCI research. Previous research on MOGs has demonstrated how virtual world(s) can create shared social norms among a group of people (Boellstorff, 2008; Nardi, 2010; Zhong, 2011; Kou and Gui, 2014; Yee, 2006) along with fostering meaningful social relationships and social support (Tong et al., 2021; Freeman et al., 2016; Lei et al., 2024). Existing work in this area lays a strong theoretical foundation for the development and study of how game communities as a whole contribute to collective in-game objectives, the empirical and knowledge contribution on how social cooperation is orchestrated at the level of individual interactions in order to address personal players’ needs is limited. Contrary to MOGs that stress sociality and cooperation (Yee, 2006), players’ experiences in Otome games are personal and customized. Accordingly, cos-commission – as a player-motivated small practice – is a useful lens through which to consider people working together to make sense of the game in a game-oriented single-player environment.
Based on semi-structured interviews with 15 players within the cos-commission ecosystem, we find that this practice allows players to co-create personal intimate experiences and to recover from negative emotions. Cos-commissions offer freedom through crafted encounters. But risks such as post-commission ”withdrawal symptoms” and blurred barriers between performative roles and actual emotions demonstrate a need for prudent supports. Though digital media technologies serve hybrid interaction in cos-commission, as is, the current designs have no means to handle these risks. Our findings support features on the platform that enable creative expression and restrict exposure, to allow gaming subcultures to formally and safely congregate.
2. Background and Related Work
Otome game, sometimes contracted to Otoge, is a story-based video game that is targeted towards women (Tian, 2022). The first commercial Otome game, Angelique, appeared in the early 1990s in Japan (Kim, 2009). As China’s mobile gaming market exploded during the 2010s, this type of game expanded by leaps and bounds and amassed a huge number of fans. Several successful mobile otome games also launched in this period, such as Mr. Love: Queen’s Choice [^1] https://evol.papegames.cn/ [Last Accessed: Nov 30, 2024], Light and Night [^2] https://love.qq.com/ [Last Accessed: Nov 30, 2024], Tears of Themis [^3] https://tot.hoyoverse.com/ [Last Accessed: Nov 30, 2024], and Love and Deepspace [^4] https://loveanddeepspace.infoldgames.com/ [Last Accessed: Nov 30, 2024]. With the development of Otome games, offline events including fan meetings and cos-commissions have become common.
In sociology, subcultures can be described as social groups that diverge from the mainstream through unique practices, signs, and shared identities (Haenfler, 2014). In HCI, subcultures have been examined for their distinct communities in relation to technology use, social norms for participation, and technology generation (Tanenbaum and Tanenbaum, 2018; Su et al., 2019; Sher and Su, 2019). The post-subcultural theory also contends that subcultures are more fragmented and more fluid and hybrid in the present, with variable practices and participants. Cosplay is a good example of this hybridization, where individual identity and collective identity are fused together through the performance of using inputs from anime, comics, games, and novels (ACGN) (Rahman et al., 2012; Ito et al., 2012).
Within the context of the purpose of cos-commission, ”closeness” is complex and consists of emotional affinity, mutual comprehension, and shared experience (Bardzell et al., 2014). Members of the Otome game communities develop intimacy in all sorts of mediated social ways online – from exchanging photos to joint participation in tasks or shared discussions. The advantage of online intimacy over offline intimacy is multifold. There is evidence that online intimacy is associated with higher levels of social support for minority members who may experience greater difficulties in creating close relationships offline (Freeman et al., 2016). In terms of accessibility, digital systems have been noticed for their affordances for intimacy, for example, anonymity, decentralization, and shared goals (Freeman et al., 2016; Li et al., 2023; Maloney and Freeman, 2020; Boellstorff, 2008). Yet little is known about the hybrid settings (combining online with offline experiences) in terms of how the co-creation of intimate experiences takes place. Cos-commission as a node between online subcultural community practices and the co-creation of intimacy offers a useful framing to start asking about this.
3. Methodology
Table 1. Demographic Information of Interviewees
ID Age Gender Education Role in cos-commission Experience
P1 21 Female Bachelor’s Cosplayer 12 months
P2 22 Female Bachelor’s Customer 24 months
P3 18 Female High School Customer 10 months
P4 22 Female Bachelor’s Cosplayer 12 months
P5 20 Female Bachelor’s Both 6 months
P6 21 Male Bachelor’s Cosplayer 12 months
P7 19 Female Bachelor’s Both 13 months
P8 20 Female Bachelor’s Cosplayer 24 months
P9 19 Female Bachelor’s Customer 13 months
P10 21 Female Bachelor’s Both 14 months
P11 Anonymous Anonymous Anonymous Anonymous Anonymous
P12 21 Female Bachelor’s Customer 10 months
P13 20 Female Bachelor’s Both 13 months
P14 22 Female Bachelor’s Both 12 months
P15 22 Female Bachelor’s Both 15 months
To understand the motivations, interactions, and challenges experienced by both customers and cosplayers in the cos-commission ecosystem, as well as their expectations for future practice, we conducted semi-structured interviews with 15 participants in China who had participated in cos-commission either as cosplayers or customers (14 females and 1 male). The demographic information of our participants is detailed in Table 1. Participants were recruited via snowball sampling. Prior to the interviews, they were briefed on the study’s purpose and background. With their consent, all interviews were recorded and transcribed verbatim in Chinese. Interview questions covered participants’ past experiences with cos-commission, their motivations, preparations for commissions, and reflections on the experience. Follow-up questions included challenges encountered in cos-commission, potential improvements, and participants’ perspectives on the future of this practice. Interviews averaged about 60 minutes, and each participant was compensated 60 CNY. Two native Mandarin-speaking authors independently coded 20% of the transcripts and then met to establish a unified codebook. The codes were iteratively organized into a thematic hierarchy (Clarke and Braun, 2017)and resulted in a consensus on the motivations, practices, challenges, and safety/privacy concerns of participants, as detailed in the following section.
4. Findings
In this section, we present the role-specific motivations, interaction dynamics, and risks associated with cos-commissions.
4.1. Role-Specific Motivations
Our preliminary findings reveal that motivations vary based on participants’ roles in the cos-commission ecosystem and their past experiences. There are also some shared motivations, such as social and emotional needs. Specifically, cosplayers pay more attention to curating identity through performative labor, while customers focus more on fulfilling their social and emotional needs.
4.1.1. Cosplayers
Although some cosplayers describe cos-commission as a part-time job, most do not expect considerable financial rewards from this activity. Many expressed ethical considerations in pricing, recognizing that most customers are students with limited income. For instance, P1 mentioned, ”Typically, these young girls (customers) don’t have much money, so I won’t set my prices too high.” For these cosplayers, the price is perceived as compensation for their time and preparation of character costumes and make-up.
Rather than financial reward, many cosplayers are driven by curiosity and a desire to pursue personal interests when accepting commissions. For example, P7, an experienced cosplayer, shared, ”My first encounter with cos-commission was at a Comic Convention, where someone saw me cosplaying a character she liked and asked if I took commission requests. Since I am also an Otome game player, I wanted to try it out. That’s how I started accepting commissions on different platforms.” This interest-driven motivation is further reflected in cosplayers’ preference for portraying characters they are familiar with. As P7 described, ”If you go on a date with someone unfamiliar with the character, you’ll get a normal dating experience, but it won’t evoke the same depth of feeling as it would with someone who truly understands the character.” For these cosplayers, monetization is seen as an additional benefit—i.e., a reward for their skill and passion—rather than a primary goal.
4.1.2. Customers
Customer motivations can be quite complex and there are many influences. Overall, the common motivators are divided into two categories: social motivators and emotional motivators.
Social Motivator
The social motivator consumers have for cos-commission is an example of the interplay between cos-commission activities and social impulses at the level of consumer practices, popular culture, and social life. For example, P2 highlighted the experience of attending conventions and cos-commissioning: ”Attending conventions with friends has been a primary motivation for me. As Otome games grew in popularity, so did conventions and cos-commission. At one convention, I saw a single customer hire six cosplayers simultaneously for group photos. That’s when I first became aware of it.” Offline events present these customers with a wealth of opportunities to grab onto Otome game cultures and at the same time support them in keeping up relationships with their friends and other game fans. Likewise, P3 described her first exposure to cos-commission as serendipity when she stumbled upon videos on social media, ”I found other people’s cos-commission videos and their sweet interaction. That’s when I figured I would give it a shot myself. I wanted to make some memories.” The increase of cos-commission coincides with the rise of Otome and Otome-related conventions, indicating further growth within the niche space of this subculture and approachability to some of the participators’ desires.
Emotional Motivator
Many customers are viewing cos-commission as a continuation of the in-game emotional satisfaction. P2 told us: ”He (the virtual character) is so perfect. Over time, I developed a desire to touch him, even though it was impossible. I want to hug him, hold his hands, and wish I could meet him in real life.” Through cos-commission, they form an opportunity to have physiological or closer communications that virtual worlds can not offer.
Others battling with mental health find cos-commission a source of emotional recovery and support otherwise missing from their lives. For example, P9 described cos-commission as the road to recovery: ”I wish I could get someone who resembles my lover to play him in real life. It would be a reward for me.” Another participant, P2, in contrast, took cos-commission to be an act of introspection: ”I want to verify through cos-commission that I have the ability to love and communicate.” Despite that there is no concrete interaction rule, the positive impact of cos-commission to mental health should not be ignored. Most of the customers who have extensive experience in Otome games or cos-commissions have their own ”dream lover”-like game characters. Cos-commission allows them to have physical contact with their ”dream lovers” as part of an attempt to turn digital connections into real ones.
4.2. Progressive Intimate Experience
In this subsection, we discuss a collective intimacy-building process that evolved in the cos-commission practice.
4.2.1. Pre-commission
Selective matching and boundary negotiation are commonly seen in the pre-commission phase. Before the commission took place, customers chose cosplayers with several criteria. Although the details are different, there are some shared concerns. Many customers and cosplayers first go through a short list of questions, including ones about physical measurements/BMI stats, etc., as well as standard queries on height, weight, nickname, level of intimacy that you are comfortable with, and relationship status. P14 considered the impression she had of the character to identify which cosplayers are more likely to be chosen: ”The appearance of the cosplayer needs to be consistent with the impression I have of the character.” Physical appearance is one of the most cited reasons when selecting a cosplayer.
4.2.2. During the Cos-commission
During a typical cos-commission, there will also be some kind of relationship progression, as well as planned interactions. Date venues and activities are usually decided by the customer. In the case that the customer cannot or do not want to make that decision, it is the cosplayer who will do so instead.
Several customers told us it feels like the dating gets better as the day goes on, that they start a little bit awkward but have a close connection by evening. P3 talked about these new roles in the form of nicknames: ”In the beginning, I used to call her ’teacher’ where we didn’t know each other and she used to call me ’sponsor’. As we got closer, I used more endearing, couple-style names or cute nicknames (e.g., ”little xxx”).” This change reflects an intimacy evolution from established comfort and trust.
4.2.3. Post-commission
After the commission, participants often have some rituals to remember and keep emotional attachments at bay. Many are showered with detailed feedback and gifts, including thank-you letters painstakingly crafted by hand or homemade mementos of gratitude. We have also observed that there is a lot of reported long-term emotional fulfillment from many of these exchanges. According to P4, ”After the date, the cosplayer said something very encouraging. If someone else had said it, I might not have listened, but because it was from my cosplayer, I did. … We promised to meet again one day. When I see the photos or other mementos from that day, I remember the promise, which gives me strength to keep going.” This long-term impact is counter-intuitive to the temporary nature of cos-commission and is evidence of the power of offline exchanges expanding the benefit of a virtual world.
4.3. Risks in Hybrid Intimacy Construction
While cos-commission offers positive experiences, there are still some underlying risks. In this subsection, we describe some common issues—withdrawal symptoms and blurred boundaries.
4.3.1. Withdrawal Symptoms
From our interviews, we noticed that almost all customers reported experiencing varying degrees of what they called ”withdrawal symptoms” afterward. This term is used to describe states of emotional confusion and difficulty readjusting to everyday life at the end of the commission. P3 likened cos-commission to a dream: ”There is a feeling that what I experienced that day was like a dream, much like Cinderella’s crystal shoes gave you a beautiful dream. When the day ends, you have to go back. It’s very painful.” While several methods have been adopted to minimize participants’ emotional attachment, as discussed in Section 4.2.3, the methods still may not be sufficient to allow participants to fully cancel out the emotional impact of cos-commissions. As a result, long-term emotional support is weakened by the feeling of loss and contrast.
Immersive environments have the ability to evoke deeper emotions and empathy (Liu et al., 2022, 2024; You et al., 2023). However, in cos-commission, unlike formal drama therapy (Landy, 1994), there is a lack of an exit mechanism in cos-commission, and the participants are not well-trained. Consequently, the deep, extensive, and dyadic interactions within a short period cause difficulties for participants in digesting and processing their emotions.
4.3.2. Blurred Boundaries between Character Embodiment and Selfhood
Currently, there is no structured guideline for cos-commission. Cosplayers typically self-learn the personality of characters by watching anime, manga, and videos to provide an immersive and high-quality service. However, unlike other forms of paid companionship—such as sex workers, game companions (Shen et al., 2021), or host/hostess clubs in Japan (Takeyama, 2020)—cos-commission participants do not share a clear, unified understanding of their short-term relationship, which may lead to unexpected and unavoidable emotional transference. As such, it often becomes difficult for participants to distinguish whether their feelings are toward the character or the cosplayer themselves, leaving a paradoxical situation where the performance ends but the emotional immersion does not.
4.4. Platform-Related Vulnerabilities
Digital platforms, as illustrated in Figure 1, are the main venues for organizing cos-commissions. However, they also present some vulnerabilities in the context of cos-commission.
4.4.1. Unregulated Information Exchange and Unwanted Contact
On different platforms, participants have a variety of options to customize their media preferences. Video-sharing websites, such as BiliBili  [^5] BiliBili: A Chinese online video-sharing platform. https://bilibili.com/ [Last Accessed: May 17, 2025], provide open venues for users to demonstrate and disseminate positive experiences, and Xiaohongshu  [^6] Xiaohongshu: Known in English as RedNote, a Chinese social networking and e-commerce platform. https://xiaohongshu.com/ [Last Accessed: May 17, 2025] (RedNote), with a fusion of graphic and textual information, enables customers to search for cosplayers, comment on reviews and service authenticity. Douyin  [^7] Douyin: A short-video platform in mainland China and Hong Kong. While TikTok and Douyin share a similar interface, they operate separately. https://douyin.com/ [Last Accessed: May 17, 2025], which is called TikTok outside mainland China and Hong Kong, specializes in short videos, allowing customers to verify the skills of cosplayers via their previous performances. A second-hand trading platform, Xianyu  [^8] Xianyu: An online second-hand trading platform. https://goofish.com/ [Last Accessed: May 17, 2025], simplifies the cos-commission process with quick communication and payments and has many who are convenience-minded users. By using these various platforms, customers can find their preferred cosplayers and improve their cos-commission experiences with easy, safe, and high-quality transactions.
4.4.2. Privacy and Safety Concerns
Although the benefits, there are also safety and privacy concerns over digital platforms. Professional cosplayers have raised apprehensions about being harassed or subjected to entanglements from a customer with whom they have built a relationship after a commission ends. For example, P10 reported that someone he knows achieved this: ”My friend posted a photo with her location on her birthday, and a customer followed the location to find her. She didn’t think the customer would, so she didn’t worry. … The customer arrived at her birthday party uninvited and insisted she drive him/her home, causing trouble for her and her girlfriend.” This is a case in point for a lack of anti-harassment and fair-use policies for paid companionship in real-world settings. Due to the feature of in-depth, dyadic communication and information exchange of cos-commission, cosplayers, and customers are pushed to a higher risk in Personally Identifiable Information (PII) disclosure.
Female cosplayers also feared uncomfortable or mistaken encounters with male customers. Some commented that men are in general more physically powerful than women, and therefore if an accident were to happen they would have a harder time defending themselves. In cos-commission, if intimate interactions are to take place, participants often move to private spaces. Since there is no aid to be had from others, their anxiety level will increase even more.
5. Conclusion and Future work
In this preliminary study, we explored the motivations, experiences, and issues of the cos-commission participants within the Otome game community in China. Our results shows that cos-commission is a hybrid space in which the participants co-create meaningful connections. It is also a medium that allows its participants to rediscover themselves as well as find emotional recovery. However, the loose line between the performing roles and real intimacy may result in some uncomfortable effects, such as harmful behaviors driven by attachment and ”withdrawal symptoms” after performing.
Our research also reveals safety and privacy issues associated with the existing cos-commission ecosystem on digital platforms. Vulnerabilities, such as unwanted post-commission entanglements when PII (e.g., a phone number, or geolocation data) gets exposed or gender inequality among female cosplayers reflect system-level breakdowns in facilitating cross-platform intimate hybrid interactions. Strong safeguards, such as chat encryption, multi-factor identity verification, and adjustable privacy filters, are required to be put in place by the platforms. These could help to create safer spaces while maintaining the independent creativity of cos-commission culture.
Through our study of subculture communities as a case-in-point, we hope to contribute to the design of virtual world interaction at large. It would be interesting for future research to consider the longitudinal dynamics of cos-commission participation, including the transformation of emotional investments across multiple visits. Together with Otome game communities, we hope to co-design socio-technical infrastructures that align immersive intimacy with appropriate safeguards. If we (re)valorized emotional satisfaction alongside participant welfare, we might learn from subcultures how to construct a better digital world.